# Zhaoxiang Zhang9, Yuanxing Zhang10, Ge Zhang3,7, Ke Xu1, Jie Fu11*, Wenhao Huang2,3

### v3

@@ 2.3 Supervised Fine-Tuning
As shown in Table 9, our model undergoes comprehensive and systematic supervised fine-tuning (SFT) with 16 different tasks and 34 diverse open-source datasets. The chat template used for SFT is the same as that used for Yi-6B-Chat (AI et al., 2024), and only the assistant responses are supervised. We refer the reader to Appendix C for more details about the hyperparameters and prompt templates.
@@ 3 Experiments
In this section, we present our quantitative evaluation results across various domains: image-related tasks (§3.1), speech-related tasks (§3.2), and video-related tasks (§3.3). Due to the lack of benchmarks for several advanced and emergent abilities of any-to-any multimodal LLMs, we also provide numerous qualitative demonstrations (§3.5) to demonstrate these capabilities. We refer the reader to Appendix D for more details, including the decoding hyperparameters and prompt templates.
@@ 3.1 Image-Related Tasks
Table 3: Experimental results for image understanding abilities. “Imagen” denotes whether the model is capable of generating images. “Speech” denotes whether the model supports speech modality. “I” denotes the instruction tuned version. The metrics used are CIDEr for COCO, MCQ accuracy for the SEED Bench, and VQA accuracy for the other tasks, following the standard procedures. In all cases, higher scores indicate better performance.
Models Imagen Speech COCO VQAv2 OKVQA VizWiz SEED Bench
@@
We compare our models with Emu (Sun et al., 2023c), SEED-LLaMA (Ge et al., 2023b), AnyGPT (Zhan et al., 2024), Flamingo (Alayrac et al., 2022), Kosmos-1 (Huang et al., 2023), MetaLM (Hao et al., 2022), IDEFICS (Laurençon et al., 2023), CM3Leon (Yu et al., 2023), InstructBLIP (Dai et al., 2023), Qwen-VL-Chat (Bai et al., 2023), and LLaVA 1.5 (Liu et al., 2023a). We evaluate our models in diverse tasks, including: (1) image captioning on MS-COCO (Lin et al., 2014) Karpathy test split with CIDEr score (Vedantam et al., 2014) as the metric, (2) three visual question-answering benchmarks, i.e., VQAv2 (Goyal et al., 2016) (test-dev split), OK-VQA (Marino et al., 2019) (val split), and VizWiz (Gurari et al., 2018), with VQA accuracy as the metric, and (3) SEED-Bench (Li et al., 2023a), a comprehensive visual question-answering benchmark including 9 dimensions with MCQ accuracy as the metric. The scores for all baselines are copied from their reports. As shown in Table 3, our MIO-Instruct is ranked in the top group among all baselines, demonstrating its competitive image understanding performance. Although SEED-LLaMA achieved better scores compared to our model, we additionally support the speech modality. It is also noteworthy that MIO, with a size of approximately 7 billion parameters, outperforms several larger models such as Emu-14B and even IDEFICS-80B.
Table 4: Image generation evaluation by CLIP-I score. “I” denotes the instruction tuned version. Higher values are better.
Models MS-COCO Flickr30K
@@
We compare our models with Emu (Sun et al., 2023c), SEED-LLaMA (Ge et al., 2023b), GILL (Koh et al., 2023), and AnyGPT (Zhan et al., 2024) for image generation. We use two benchmarks, i.e., MS-COCO (Lin et al., 2014) Karpathy test split and Flickr30K (Plummer et al., 2015). Following GILL (Koh et al., 2023) and SEED-LLaMA (Ge et al., 2023b), we use CLIP-I as the metric that evaluates the similarity between the generated images and the ground-truth images with the image encoder in CLIP (Radford et al., 2021). As shown in Table 4 and Table 12 the pre-trained model and instruction-tuned model of MIO both have competitive image generation capabilities. Note that beyond single image generation abilities, our model can also exhibit multi-image generation capabilities such as generating visual stories, image sequences, and even visual thoughts as illustrated in §3.5.
@@ 3.2 Speech-Related Tasks
Table 5: Speech ability evaluation. “WER” denotes word error rate. Lower values are better.
Models ASR WER Models TTS WER
@@
As shown in Table 3.2, our models exhibit speech performance comparable to the speech-specific baselines and outperform the AnyGPT baseline. It is important to note that although AnyGPT is capable of generating content tokens for speech, it lacks the ability to generate timbre tokens, which necessitates the use of an additional voice cloning model. In contrast, our models generate both content and timbre tokens, making the TTS tasks more challenging for our models compared to AnyGPT. Nonetheless, after instruction tuning, our model still achieves better TTS performance. More evaluations of the TTS and Speech-to-Speech generation performance are provided in Appendix E.3 and E.2.
Table 6: Video understanding evaluation using top-1 accuracy for both benchmarks. “I” denotes the instruction-tuned version.
@@ 3.3 Video-Related Tasks
We compare MIO with Flamingo (Alayrac et al., 2022), BLIP-2 (Li et al., 2023b), InstructBLIP (Dai et al., 2023), Emu (Sun et al., 2023c), and SEED-LLaMA (Ge et al., 2023b) for video understanding. The models are evaluated on the MSVDQA (Chen & Dolan, 2011a) and MSRVTT-QA (Xu et al., 2017). The results are presented in Table 6. Our model achieves the highest scores compared to all baselines. Due to the lack of video (frame sequence) generation benchmarks in our setting, we provide video generation examples in §3.5. These results demonstrate the superior performance of our models in both video understanding and video generation.
Table 7: Language-only evaluation. “I” denotes the instruction-tuned version.
Models MMLU
@@
Table 8: Results for trimodal comprehension (text, image, and speech).
Models OmniBench
@@ 3.4 Language-only Tasks
We evaluate our models on MMLU (Hendrycks et al., 2021). The baselines are two LLaMA variants (Touvron et al., 2023a; b), the instruction-tuned SEED-LLaMA (Ge et al., 2023b), and AnyGPT (Zhan et al., 2024). For the MMLU benchmark, we conduct zero-shot evaluation experiments using the official evaluation code. The experimental results are shown in Table 8. We can observe that our models have superior language-only performance compared with all any-to-any MM-LLM baselines and even surpass LLaMA-1-7B-Base, an advanced pure language model.
3.5 Demonstrations
We illustrate the basic and advanced abilities of MIO in Figure 5 and 4. The basic abilities of MIO involve image understanding and generation, video understanding and generation, ASR, and TTS. The advanced abilities of MIO are based on its any-to-any and multimodal interleaved sequence generation features. These abilities involve visual storytelling (i.e., interleaved video-text generation), chain of visual thought, speech-in-speech-out, instructional image editing, visual guideline generation, etc. We refer the readers to Appendix E.5 for more demonstrations including multimodal chain of thought and multimodal in-context learning.
@@ 3.6 Ablation Studies
We evaluate our model using the OmniBench (Li et al., 2024d), which incorporates text, image, and speech modalities as inputs, requiring the model to choose one of four options as the correct answer to determine accuracy. Although MIO acquires its multimodal understanding capabilities through dual-modal training, the evaluation results in Table 8 indicate that MIO also exhibits superior trimodal comprehension abilities.
Effect of Different Image Tokenizers.
The image tokenizer has a significant impact on image modality alignment. In Figure 2, we compare the image generation performance under a controlled setting after training for solely 3K steps in stage 1, using various image tokenizers. The image tokenizers used for comparison include a VQGAN (Esser et al., 2020) with a vocabulary size of 1024 and a compression rate of 16 (VQGAN-1024), as well as the VQGAN-Gumbel with a vocabulary size of 8192 (VQGAN-8192) [^5] https://github.com/CompVis/taming-transformers. Our results indicate that the SEED-Tokenizer, which captures more semantic and higher-level image information, exhibits faster convergence. In contrast, both VQGAN tokenizers show slower convergence due to their lower-level image information.
@@ 4 Related Works
4.1 Multimodal LLMs
With the rapid success of Large Language Models (LLMs), current multimodal LLMs (MM-LLMs) are typically built on a pre-trained LLM backbone and are endowed with the ability to understand multiple modalities (Li et al., 2019; Lu et al., 2019; Kim et al., 2021; Zeng et al., 2022; Zhou et al., 2022; Wang et al., 2023b; 2024e). Generally, these MM-LLMs align the representations of images obtained from visual encoders with the text embedding space, thereby leveraging the powerful capabilities of the foundational models. For example, BLIP-2 (Li et al., 2023b) uses CLIP-ViT (Radford et al., 2021) to extract high-level features from images and then employs a Q-Former to compress the number of image tokens and further align image tokens with the LLM embeddings. In contrast, LLaVA (Liu et al., 2023b; Li et al., 2024a) utilizes a simple linear projection or MLP as the connector between the image encoder and the LLM backbone. These models demonstrate strong multimodal understanding abilities, achieving significant progress in tasks such as visual question answering, visual commonsense reasoning, chart understanding, etc.
Additionally, beyond images, other MM-LLMs have also focused on modalities such as speech and video. For instance, LLaSM (Shu et al., 2023) and InternVideo (Wang et al., 2022; 2024c) are MM-LLMs designed for speech and video understanding, respectively. These models adopt a similar architectural design to BLIP-2 or LLaVA but redesign modality-specific encoders.
Recently, increasing attention has been paid to unifying multiple modalities within a single MM-LLM. For example, ImageBind (Girdhar et al., 2023) develops encoders suited for multiple modalities such as images, videos, audio, heat maps, among others, while OmniBind (Wang et al., 2024d) trains an omni-representation model by aligning encoders across four modalities: audio, language, images, and 3D objects. OmniBench (Li et al., 2024d) is proposed to evaluate the models’ abilities for visual, acoustic, and textual understanding.
However, these models focus primarily on multimodal understanding and often overlook the important aspect of multimodal generation.
4.2 Any-to-Any MM-LLMs
To enable multimodal generation in MM-LLMs, a straightforward approach is to allow these models to call external multimodal generation tools, such as Stable Diffusion (Rombach et al., 2022) or text-to-speech (TTS) tools (Shen et al., 2023; Li et al., 2024c; OpenAI et al., 2023). However, as highlighted in the Gemini technical report (Team et al., 2023), relying on an intermediate natural language interface can limit the model’s ability to express images. If a model cannot natively output images, it will not be able to generate images with prompts of interleaved sequences of image and text. This claim is in line with our distinction between descriptive image generation and context-aware image generation, as discussed in §2.2.
As a result, recent works focus on the unification of multimodal understanding and generation in a single model (i.e., any-to-any MM-LLMs), enabling the generation of multimodal tokens without natural language as an interface. These models typically follow different approaches, depending on how images are represented in both input and output sides. For example, the Discrete-In-Discrete-Out (DIDO) approach has been explored in works such as SEED-LLaMA (Ge et al., 2023b), AnyGPT (Zhan et al., 2024), and Chameleon (Team, 2024). Continuous-In-Discrete-Out (CIDO) methods have been implemented in models like DaVinCi (Diao et al., 2023), Gemini (Team et al., 2023), and Unified-IO 2 (Lu et al., 2023). The Continuous-In-Continuous-Out (CICO) approach is used in models such as Emu (Sun et al., 2023c; a), and DreamLLM (Dong et al., 2023). Another approach, the integration of autoregression and diffusion (AR + Diff), can be seen in models like Transfusion (Zhou et al., 2024), Show-o (Xie et al., 2024), and Li et al. (2024b)’s.
@@ 4.2 Any-to-Any MM-LLMs
However, these models face specific limitations. DreamLLM (CICI,  Dong et al. (2023)) and CIDO models suffer from inconsistencies between input and output forms for multimodal data, making it difficult for them to natively support the generation of interleaved multimodal sequences where an image functions in a coupled way as both input and output. Emu2 (CICO,  Sun et al. (2023a)) struggles with the challenges of the mean square error (MSE) loss used for training continuous output representations, as well as with the uni-modal assumption of the Gaussian distribution in the MSE loss. Transfusion (AR + Diff,  Zhou et al. (2024)) applies noise to images from the input side to support multimodal generation with diffusion modeling, and relies on VAE (Kingma & Welling, 2013) features rather than CLIP (Radford et al., 2021) features for denoising, which largely trade off the multimodal understanding abilities.
To mitigate these issues, we adopt the DIDO approach. A comprehensive comparison of our models with other any-to-any MM-LLMs is presented in Table 1.
@@ Appendix A Pre-training Data
    Image-text interleaved data: OBELICS (Laurençon et al., 2023), MMC4-core-ff (Zhu et al., 2023).
@@ Appendix B Pre-training Details
We enable Flash Attention (Dao et al., 2022; Dao, 2023) during pre-training. Gradient clipping is set to 1.0 for all stages. The maximum sequence length for training is 2800 tokens. We use a cosine learning rate scheduler with a peak learning rate of 3e-5 and a warmup ratio of 0.03. The optimizer used is AdamW (Loshchilov & Hutter, 2017).

### v4

@@ 2.3 Supervised Fine-Tuning
As shown in Table 10, our model undergoes comprehensive and systematic supervised fine-tuning (SFT) with 16 different tasks and 34 diverse open-source datasets. The chat template used for SFT is the same as that used for Yi-6B-Chat (AI et al., 2024), and only the assistant responses are supervised. We refer the reader to Appendix C for details about the hyperparameters and prompt templates.
Model Overall Understanding Generation Unify
Gemini2.0-flash-exp 45.6 65.2 29.8 40.7
MIO-Instruct 37.2 41.5 53.5 16.6
SEED-LLaMA 28.5 39.5 23.5 22.3
Anole 18.6 13.6 20.0 22.3
VILA-U 18.6 40.0 15.8 -
Janus-Pro 18.1 48.4 5.9 -
Janus-Flow 16.3 43.4 5.5 -
Emu3 13.8 33.2 8.2 -
Show-o 12.7 31.0 7.3 -
Table 2: Scores on MME-Unify (Xie et al., 2025). “-” indicates that the model lacks the capability to complete the specified task. Understanding tasks involve single image perception & understanding, multiple & interleaved image-text understanding, and video perception & understanding. Generation tasks include conditional image-to-video generation, fine-grained image reconstruction, text-guided image editing, text-to-image generation, text-to-video generation, and video prediction. Unified tasks encompass image editing and explaining, common sense question answering, auxiliary lines, SpotDiff, and visual chain-of-thought reasoning. Arrange from top to bottom based on the overall score from highest to lowest.
@@ 3 Experiments
In this section, we first report our scores on MME-Unify (Xie et al., 2025), a widely used leaderboard for evaluating the multimodal performance of any-to-any MLLMs. Subsequently, we present our quantitative evaluation across various domains: image-related tasks (§3.2), speech-related tasks (§3.3), and video-related tasks (§3.4). Due to the lack of benchmarks for several advanced and emergent abilities of any-to-any multimodal LLMs, we provide qualitative demonstrations (§E.5) demonstrating these capabilities. The decoding hyperparameters and prompt templates are shown in Appendix D.
3.1 Scores on MME-Unify
We report our scores on a well-recognized third-party leaderboard for unified models, i.e., MME-Unify (Xie et al., 2025), in Table  2. According to the Overall metric on this leaderboard, our MIO model surpasses models such as Janus-Pro (Chen et al., 2025a) [2] and achieves multimodal unified modeling capabilities second only to Gemini-2.0-Flash, on the condition that MIO covers a broader range of tasks compared to Gemini-2.0-Flash, demonstrating MIO’s top-tier multimodal unified modeling capabilities. Furthermore, MIO exhibits impressive multimodal understanding capabilities, ranking highly among unified MLLMs according to the Understanding metric, and demonstrates exceptionally leading performance in generation tasks, as evidenced by the Generation metric. Additionally, MIO extends its capabilities to include speech support, a feature not offered by many competing models. However, we observe that MIO demonstrates limited performance on the unified tasks. This is primarily because the unified tasks in MME-Unify involve single-choice questions with images as options, a type of data that MIO’s training lacks.
@@ 3.2 Image-Related Tasks
Models Imagen Speech COCO($\uparrow$) VQAv2($\uparrow$) OKVQA($\uparrow$) VizWiz($\uparrow$) SEED Bench($\uparrow$)
@@
Table 3: Experimental results for image understanding abilities. “Imagen” denotes whether the model is capable of generating images. “Speech” denotes whether the model supports speech modality. “I” denotes the instruction tuned version. The metrics used are CIDEr for COCO, MCQ accuracy for the SEED Bench, and VQA accuracy for the other tasks, following the standard procedures.
@@
We compare our models with Emu (Sun et al., 2023c), SEED-LLaMA (Ge et al., 2023b), AnyGPT (Zhan et al., 2024), Flamingo (Alayrac et al., 2022), Kosmos-1 (Huang et al., 2023), MetaLM (Hao et al., 2022), IDEFICS (Laurençon et al., 2023), CM3Leon (Yu et al., 2023), and InstructBLIP (Dai et al., 2023). We evaluate our models in diverse tasks: (1) image captioning on MS-COCO (Lin et al., 2014) Karpathy test split with CIDEr score (Vedantam et al., 2014) as the metric, (2) three visual question-answering benchmarks, i.e., VQAv2 (Goyal et al., 2016) (test-dev split), OK-VQA (Marino et al., 2019) (val split), and VizWiz (Gurari et al., 2018), with VQA accuracy as the metric, and (3) SEED-Bench (Li et al., 2023a), a comprehensive visual question-answering benchmark including 9 dimensions with MCQ accuracy as the metric. The scores for all baselines are copied from their reports. As shown in Table 3, our MIO-Instruct is ranked in the top group among all baselines, demonstrating its competitive image understanding performance. Although SEED-LLaMA achieved better scores, we additionally support the speech modality. It is noteworthy that MIO, with a size of approximately 7 billion parameters, outperforms several larger models such as Emu-14B and even IDEFICS-80B.
Models MS-COCO($\uparrow$) Flickr30K($\uparrow$)
@@
Table 4: Image generation evaluation by CLIP-I score.
@@
We compare our models with Emu (Sun et al., 2023c), SEED-LLaMA (Ge et al., 2023b), GILL (Koh et al., 2023), and AnyGPT (Zhan et al., 2024) for image generation. We use two benchmarks: MS-COCO (Lin et al., 2014) Karpathy test split and Flickr30K (Plummer et al., 2015). Following GILL (Koh et al., 2023) and SEED-LLaMA (Ge et al., 2023b), we use CLIP-I as the metric that evaluates the similarity between the generated and the ground-truth images with the CLIP image encoder (Radford et al., 2021). As shown in Table 4 and 13, the pre-trained and instruction-tuned model of MIO both have competitive image generation abilities. Note that beyond single image generation, our model can also exhibit multi-image generation capabilities such as generating visual stories, image sequences, and visual thoughts as illustrated in §E.5.
@@ 3.3 Speech-Related Tasks
Models ASR Models TTS
@@
Table 5: Speech ability evaluation by WER ($\downarrow$).
As shown in Table 5, our models exhibit speech performance comparable to the speech-specific baselines and outperform the AnyGPT baseline. It is important to note that although AnyGPT is capable of generating content tokens for speech, it lacks the ability to generate timbre tokens, which necessitates the use of an additional voice cloning model. In contrast, our models generate both content and timbre tokens, making the TTS tasks more challenging for our models compared to AnyGPT. Nonetheless, after instruction tuning, our model still achieves better TTS performance. More evaluations of the TTS and Speech-to-Speech generation performance are provided in Appendix E.3 and E.2.
@@
Table 6: Video understanding evaluation using accuracy.
@@ 3.4 Video-Related Tasks
We compare MIO with Flamingo (Alayrac et al., 2022), BLIP-2 (Li et al., 2023b), InstructBLIP (Dai et al., 2023), Emu (Sun et al., 2023c), and SEED-LLaMA (Ge et al., 2023b) for video understanding. The models are evaluated on MSVDQA (Chen and Dolan, 2011a) and MSRVTT-QA (Xu et al., 2017). The results are presented in Table 6. Our model achieves highest scores compared to all baselines. Due to the lack of video generation benchmarks in our setting, we provide examples in §E.5. These results demonstrate superior performance of our models in both video understanding and generation.
Models MMLU($\uparrow$)
@@
Table 7: Language-only evaluation.
Models OmniBench($\uparrow$)
@@ 3.5 Language-only Tasks
We evaluate our models on MMLU (Hendrycks et al., 2021). The baselines are two LLaMA variants (Touvron et al., 2023a, b), the instruction-tuned SEED-LLaMA (Ge et al., 2023b), and AnyGPT (Zhan et al., 2024). For the MMLU benchmark, we conduct zero-shot evaluation experiments using the official evaluation code. The experimental results are shown in Table 7. We can observe that our models have superior language-only performance compared with all any-to-any MM-LLM baselines and even surpass LLaMA-1-7B-Base, an advanced pure language model.
@@ 3.6 Ablation Studies
We evaluate our model using OmniBench (Li et al., 2024c), which incorporates text, image, and speech modalities as inputs, requiring the model to choose one of four options as the correct answer to determine accuracy. Although MIO acquires its multimodal understanding abilities via dual-modal training, the results in Table 8 indicate that MIO exhibits superior trimodal comprehension abilities.
Please refer to Appendix E.6 for more ablations including the effect of different image tokenizers.
@@ 4 Related Works
With the success of LLMs, MM-LLMs have emerged, extending LLMs to handle images, speech, and video (Liu et al., 2023; Li et al., 2024a, 2023b; Bai et al., 2023; Team, 2025; OpenAI, 2023; Shi et al., 2025; Chen et al., 2025b). These models typically align image features with text embeddings. For instance, BLIP-2 (Li et al., 2023b) uses CLIP-ViT and a Q-Former for alignment, while LLaVA (Liu et al., 2023; Li et al., 2024a) uses linear projections. These models perform well in visual question answering and commonsense reasoning. Recent models like LLaSM (Shu et al., 2023) and Qwen2.5-VL (Team, 2025) extend to speech and video. However, most focus on understanding rather than generation, limiting their utility in fully multimodal tasks.
To address this, recent work explores any-to-any MM-LLMs that generate outputs across modalities without intermediate language. Key approaches include Discrete-In-Discrete-Out (DIDO), Continuous-In-Discrete-Out (CIDO), Continuous-In-Continuous-Out (CICO), and Autoregressive + Diffusion (AR + Diff), discussed in §1. DIDO is used in SEED-LLaMA (Ge et al., 2023b), AnyGPT (Zhan et al., 2024), and Chameleon (Team, 2024); CIDO in DaVinCi (Diao et al., 2023), Gemini (Team et al., 2023), and Unified-IO 2 (Lu et al., 2023); CICO in Emu (Sun et al., 2023c, a) and DreamLLM (Dong et al., 2023); AR + Diff in Transfusion (Zhou et al., 2024), Show-o (Xie et al., 2024), and MAR Li et al. (2024b).
@@
However, these models have limitations. DreamLLM (CICO,  Dong et al. (2023)) and CIDO models suffer from inconsistencies between input and output forms for multimodal data, making it hard to generate interleaved multimodal sequences where an image functions in a coupled way as both input and output. Emu2 (CICO,  Sun et al. (2023a)) faces challenges with MSE loss for training continuous outputs, as well as with the uni-modal assumption of the Gaussian distribution in the MSE loss. Transfusion (AR + Diff,  Zhou et al. (2024)) applies noise to images from the input side to support multimodal generation with diffusion modeling, and relies on VAE (Kingma and Welling, 2013) rather than CLIP (Radford et al., 2021) features for denoising, which largely trade off the multimodal understanding abilities. To mitigate these issues, we adopt the DIDO approach. A comprehensive comparison of our models with other any-to-any MM-LLMs is presented in Table 1.
@@ 5 Conclusion
Limitations
While MIO demonstrates advancements in multimodal understanding and generation, it has certain limitations. First, the model’s performance is constrained by the quality and diversity of the training data, particularly for speech and video modalities, where open-source datasets may not fully capture the range of real-world scenarios. Second, the computational complexity of handling four modalities simultaneously requires substantial resources, potentially limiting scalability and accessibility for smaller research groups. Third, while MIO excels in multimodal interleaved sequence generation, the evaluation of such capabilities lacks standardized benchmarks, making it challenging to quantitatively compare with other models. Finally, the model’s ability to handle extremely long-context multimodal sequences or highly specialized tasks may be limited due to the fixed context window. Future work could address these by expanding dataset diversity, optimizing computational efficiency, and developing tailored evaluation metrics for advanced multimodal tasks.
Acknowledgements
This work was supported in part by the Jiangsu Science and Technology Major Project (BG2024031) and Nanjing University AI & AI for Science Funding (2024300540). It was also supported by the Shanghai Artificial Intelligence Laboratory.
@@ Appendix A Pre-training Data
    Image-text interleaved data: OBELICS (Laurençon et al., 2023), MMC4-core-$f$ (Zhu et al., 2023).
@@ Appendix B Pre-training Details
Pre-training Stage Stage I Stage II Stage III
Language-Only RefinedWeb RefinedWeb RefinedWeb
Image-Text Inter - OBELICS, MMC4-core-ff MMC4-core-$f$
Video-Text Pair - WebVid-10M WebVid-10M
Speech-Text Pair Libriheavy Libriheavy Libriheavy
GPUs 128 A800-80GB 128 A800-80GB 8 A800-80GB
Training Steps 24,800 12,800 32,200
Batch Size 12:2:0:2 2:2:6:6 2:1:1:12
@@
We enable Flash Attention (Dao et al., 2022; Dao, 2023) during pre-training. Gradient clipping is set to 1.0 for all stages. The maximum sequence length for training is 2800 tokens. We use a cosine learning rate scheduler with a peak learning rate of 3e-5 and a warmup ratio of 0.03. The optimizer used is AdamW (Loshchilov and Hutter, 2017).